OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 67. n + (n + 2) = 67 n − (n + 1) = 67 n + n = 67 n + (n + 1) = 67

OpenStudy (anonymous):

guess and check

OpenStudy (anonymous):

i bet you could guess it on the first try

OpenStudy (anonymous):

the reason that the answer is n+(n+1)=67 is this: lets say that the two numbers are x and y so the question is asking us to find two numbers that make x+y=67 true but they also tell us that y=x+1 so if we substitute: we get x+(x+1)=67 replace, x with n and you have your answer

OpenStudy (anonymous):

let me be quiet.